OpenStudy (anonymous):

Which of the following could be used to find the height of the flagpole shown?

OpenStudy (anonymous):

x = 5 cos 40° x = 5 cos 50° x = 5 sin 40° x = 5 tan 50°

OpenStudy (ranga):

Are you sure you typed all the answer choices correctly? Let 'x' be the height of the flagpole. sin(50) = opposite / hypotenuse = x / 5 x = 5sin(50).

OpenStudy (anonymous):

thank you .

OpenStudy (ranga):

Oh, BTW, sin(50) is same as cos(40). So 5sin(50) = 5cos(40). So (A).

OpenStudy (anonymous):

oh ohkay i was just about to ask

OpenStudy (ranga):

sin(x) = cos(90-x). sin(50) = cos(90-50) = cos(40). This is true of ALL "co" trig functions: sine and "co"sine, tan and "co"tan or cot, sec and "co"sec or csc. tan(x) = cot(90-x) and cot(x) = tan(90-x), etc..